# Interpretable Fine-Tuning and Error Indication for Graph Neural Network Surrogate Models

### v3

@@ Abstract
Data-driven surrogate modeling has surged in capability in recent years with the emergence of graph neural networks (GNNs), which can operate directly on mesh-based representations of data. The goal of this work is to introduce an interpretable fine-tuning strategy for GNNs, with application to unstructured mesh-based fluid dynamics modeling. The end result is an enhanced fine-tuned model that isolates regions in physical space, corresponding to sub-graphs, that are intrinsically linked to the forecasting task while retaining the predictive capability of the baseline. These structures, identified by the fine-tuned GNNs, are adaptively produced in the forward pass and serve as explainable links between the baseline model architecture, the optimization goal, and known problem-specific physics. Additionally, through a regularization procedure, the fine-tuned GNNs can also be used to identify, during inference, graph nodes that correspond to a majority of the anticipated forecasting error, adding a novel interpretable error-tagging capability to baseline models. Demonstrations are performed using unstructured flow field data sourced from flow over a backward-facing step at high Reynolds numbers, with geometry extrapolations demonstrated for ramp and wall-mounted cube configurations.
@@ 1 Introduction
The goal of this work is to introduce additional, equally critical advantages that can be provided within the GNN framework: (1) the ability to generate interpretable latent spaces (or latent graphs) as a by-product of the network architecture through a fine-tuning procedure, and (2) the ability to use these interpretable latent graphs to adaptively identify, or tag, sub-graphs that contribute most to the prediction error (i.e., a-posteriori error tagging). To emphasize broad applicability, these goals are achieved here in the context of fine-tuning: more specifically, when given a pre-trained baseline GNN surrogate model, this work shows how appending an additional graph-based trainable module can not only enhance the interpretability properties of the baseline in various modeling tasks (with focus on mesh-based fluid flow forecasting), but also provide key modeling benefits in the form of error tagging. This strategy is illustrated in Fig. 1.
Figure 1: (Top, blue box) Schematic of baseline GNN model architecture. (Bottom, red box) Schematic of interpretable fine-tuning module. Module is implemented via skip connection between first and last message passing layers in baseline processor block. Interpretability comes from adaptive pooling layer, which generates sub-graphs that can be visualized as masked fields. Using a regularization strategy, masked fields can then be used to identify regions of forecasting approximation errors during model evaluation. Details provided in Sec. 3.1. (Bottom, green box) Error tagging procedure by means of a budget regularization strategy that encourages prediction errors to be contained within the extracted nodes in sub-graphs (the masked fields). Details provided in Sec. 3.2.
@@
The scope of this previous work [34], however, concerns unstructured flow field reconstruction (i.e., a GNN was used to learn an interpretable identity map). Instead, the objective of this work is to (a) extend this notion of interpretability into a predictive mesh-based modeling setting, with focus on forecasting unstructured turbulent fluid flows, and (b) use such interpretability enhancements to append black-box models with explainable error-tagging capability. As such, given an input graph-based representation of an unstructured flow field, the GNN in this work models the time-evolution of node features. For this forecasting task, the objective is to show how augmenting pre-trained baseline GNNs with trainable adaptive pooling modules results in an interpretable fine-tuning framework for mesh-based predictive modeling. Ultimately, the input to this procedure is an arbitrary GNN baseline model, and the output is a modified GNN that produces interpretable latent graphs tailored to different fine-tuning objectives, including error-tagging capability in the context of data-based forecasting. Although the application focus here is mesh-based fluid flow modeling, the framework is expected to extend into other applications that leverage graph neural networks. The core elements of this work are as follows:
@@
    Multiscale GNN baseline: A baseline GNN leveraging multiscale message passing (MMP) layers is trained to model the evolution of an unstructured turbulent fluid flow. Datasets to demonstrate the methods introduced in this work are sourced from unstructured fluid dynamics simulations of flow over a backward facing step using OpenFOAM, an open-source CFD library [35]. Reynolds numbers of flow trajectories range from 26000 to 46000.
@@
    Interpretable fine-tuning: The aforementioned graph pooling layer – which relies on a learnable node subsampling procedure – is attached to the baseline GNN, creating an augmented GNN architecture with enhanced interpretability properties. Fine-tuning is then accomplished by freezing the parameters of the baseline GNN and ensuring the parameters in the newly introduced module are freely trainable.
@@
    Error tagging: A regularization term, representing a mean-squared error budget, is added to the objective function during the fine-tuning process. The minimization of this term ensures that the fine-tuned GNN tags, during inference, a subset of the graph nodes that are expected to contribute most significantly to the GNN forecasting error.
@@
    Geometry extrapolation: Fine-tuned GNNs – which were trained on the BFS geometry alone – are evaluated on two additional geometries exhibiting different flow separation physics (a ramp and a wall-mounted cube case). The objective is to assess the level of consistency in identified features and error tagging capability by means of inference on unseen configurations.
@@ 2 Methods
Mesh-based data required to train the GNNs are sourced from unstructured computational fluid dynamics (CFD) simulations in a two-dimensional backward-facing step (BFS) configuration using OpenFOAM [36]. Simulations are performed for a range of Reynolds numbers (Re) based on the BFS step height to populate training and testing sets. Specifically, 10 training set trajectories were sampled in the equi-spaced Re range of [26214, 44569], and 10 testing set trajectories in the Re range of [27233, 45589]. Each trajectory is comprised of $419$ two-component velocity snapshots separated by $\Delta t=10^{-4}s$.
@@
    Absence of pressure: The Navier-Stokes equations, which are the governing equations for fluid flow, identify the role of local pressure gradients on the evolution of fluid velocity. The GNN in Eq. 2, however, operates only on velocity field data, and does not explicitly include pressure or its gradients in the evaluation of dynamics. Instead, the model is tasked to implicitly recover the effect of pressure on the flow through data observations. This is consistent with real-world scenarios in which time-resolved velocity data is accessible (e.g., through optical flow [38] or laser-based imaging [39]), but pressure field data (particularly pressure gradient data) is inaccessible or sparse.
@@ 3 Results
3.1 Fine-Tuning GNNs for Interpretability
It is assumed that a black-box graph neural network optimized for the above described forecasting task – termed the baseline GNN, ${\cal G}_{B}$ – is available. With ${\cal G}_{B}$ as the starting point, the fine-tuning procedure (to be described and demonstrated in this section) produces a new GNN – termed the fine-tuned GNN, ${\cal G}_{F}$ – that is both physically interpretable and optimized for the same modeling task as the baseline ${\cal G}_{F}$. These interpretability properties are embedded into the baseline through architectural modifications; more specifically, ${\cal G}_{F}$ is created by appending ${\cal G}_{B}$ with an interpretable fine-tuning module that does not modify the baseline parameters. The module provides interpretability through a learnable pooling operation that adaptively isolates a subset of nodes in the context of the mesh-based forecasting task (the objective function in Eq. 3). It will be shown how these nodes can be readily visualized to extract coherent physical features most relevant to this forecasting goal, all while retaining the architectural structure and predictive properties of the baseline model. This approach is illustrated in Fig. 1.
@@ 3.1.1 Baseline GNN
Message passing layers can take various forms and are the backbone of all GNN architectures, since they directly leverage the graph connectivity to learn complex functions conditioned on the arrangement of graph nodes in neighborhoods [12]. The processor in the baseline architecture in Fig. 1 is a general description, in that the fine-tuning procedure is not tied down to the type of MP layer used. Here, all MP layers correspond to multiscale message passing (MMP) layers, which leverage a series of coarse grids in the message passing operation to improve neighborhood aggregation efficiency [16, 34, 15]. Details on the MMP layer used here can be found in Appendix C.
3.1.2 Interpretable Fine-Tuning Module
The interpretable fine-tuning module is shown in Fig. 1 (bottom), and contains a separate set of layers designed to enhance interpretability properties of the black-box baseline. The module is implemented through a skip connection between the first and last MP layers in the baseline processor block, such that module parameters are informed of the action of baseline neighborhood aggregation functions during parameter optimization. There are three components to the module: an adaptive graph pooling layer, an additional MMP layer that acts on the pooled graph representation, and a parameter-free un-pooling layer.
@@ 3.1.2 Interpretable Fine-Tuning Module
As illustrated in Fig. 1, the sub-graph produced by the pooling layer can be readily visualized in a mesh-based setting, since the extracted nodes correspond to locations in physical space. If the predictive capability of the baseline is withheld when optimizing the parameters in the interpretable module, nodes identified in the sub-graphs serve as an accessible link to the forecasting task. The discussion below shows how adaptive node sub-sampling provided by the pooling operation identifies physically coherent artifacts in the mesh-based domain, thereby connecting the data-based modeling task (the goal of the baseline) with application-oriented features of interest (the goal of the interpretable fine-tuning module).
3.1.3 Fine-tuning Procedure
Fine-tuning is accomplished by first initializing the ${\cal G}_{F}$ model parameters directly from those in the baseline ${\cal G}_{B}$, which is assumed to already have been trained. Then, the interpretable module is appended to the ${\cal G}_{F}$ architecture as per Fig. 1.
A training step for ${\cal G}_{F}$ is then carried out using the same objective function used to train the baseline ${\cal G}_{B}$. During training, the subset of parameters in ${\cal G}_{F}$ corresponding to the baseline is frozen, and the parameters corresponding to the interpretable module are freely trainable. This ensures that the fine-tuned GNN parameters include the baseline parameters as a subset, and isolate optimization focus to the interpretable module parameters during training; in other words, if the baseline parameters are kept frozen, the adaptive pooling layer is forced to identify sub-graphs that result in improved baseline accuracy from the perspective of the objective function being optimized (in this case, a single-step forecast). The fine-tuning procedure is shown in Fig. 3, which displays objective function optimization histories for fine-tuned GNNs trained with different values of node reduction factor $RF$.
Figure 3: Mean-squared error objective (Eq. 3) versus training epochs during fine-tuning phase.
The training histories show two key trends: (1) the baseline objective history is frozen, as the baseline model ${\cal G}_{B}$ is not being modified, and (2) the fine-tuned models achieve lower converged mean-squared errors in single-step GNN predictions than the baseline. The latter trend is consistent with the fact that additional parameters are introduced in the fine-tuning procedure (see Fig. 1). Due to the fact that a message passing layer acts on the Top-K-identified subgraphs, the converged training errors increase with decreasing RF. Higher RF translates to sub-graphs occupying a smaller percentage of original graph nodes, which in turn limits the scope of the additional message passing layer.
@@ 3.1.4 Masked Fields
The interpretability enhancement provided by the fine-tuning procedure is illustrated in Fig. 4, which contrasts the single-step prediction workflow of the baseline GNN with that of a series of fine-tuned GNNs, each having incrementally higher values of RF. The input to all models is the same instantaneous mesh-based flowfield ${\bf X}_{m}$ sourced from a testing set trajectory.
Figure 4 shows how the fine-tuning procedure adds masked fields to the model evaluation output. Simply put, a masked field is generated by visualizing the sub-graph node indices produced by the Top-K pooling operation. Since these truncated nodes coincide with physical space locations consistent with the underlying mesh, the “active” nodes selected during pooling can be extracted and visualized as a node-based categorical quantity (1 if the node was identified and 0 otherwise).
The features contained in these fields allow one to interpret the relationship between the input flow field ${\bf X}_{m}$ and the model-derived forecast $\widetilde{\bf X}_{m+1}$. In other words, through visualization of the masked fields, the fine-tuned models allow one to access regions in physical space intrinsically linked to the modeling task (single-step prediction), thereby making the model behavior directly accessible from the perspective of the objective function used during training. The masked fields in Fig. 4 show how this added capability translates to interpretable prediction unavailable in the baseline, as the masked fields interestingly isolate nontrivial, but spatially coherent, clusters of nodes (often disjoint in physical space) within the GNN forward pass. Such features can then be connected to the problem physics in an expert-guided analysis phase.
Figure 4: Comparison of GNN prediction workflows between baseline and fine-tuned GNNs. Fine-tuned workflow includes masked fields generated in forward pass. Input flowfield (top) sourced from Re=27,233 trajectory of the testing set. Bottom-most plots show baseline absolute error fields ($\widetilde{\bf X}_{m+1}-{\bf X}_{m+1}$) for streamwise (left) and vertical (right) velocity components.
@@
Figure 5: Single-step prediction outputs generated by fine-tuned GNN (RF=16) using input fields at time-ordered instants in the Re=27233 trajectory. Each frame shows predicted y-component of velocity (top) alongside masked field (bottom), with inlays showing subgraph representations.
Since the input velocities stored on the nodes are time-evolving, and the Top-K projection operation is conditioned on these input values, the regions identified in the masked fields are necessarily time-evolving as well. Consequently, the masked fields identify coherent structures that evolve in accordance with the unsteady features of the flow – in the BFS configuration, these features correspond to emerging and downstream shedding of recirculation zones, which in turn lead to a shedding cycle that affects the location at which the flow reattaches to the bottom wall. This time-dependent quality is highlighted in Fig. 5, which shows a set of time-ordered single-step predictions paired with corresponding masked fields during one BFS shedding cycle. It should be noted that the adaptive pooling operation induces a sub-graph from which the masked field is derived; since the identified regions in the masked fields are time-evolving, the adjacency matrix that characterizes this sub-graph is time-evolving as well, and remains coincident with the fixed adjacency matrix of the original mesh-based graph. The sub-graphs are provided as inlaid plots in Fig. 5; these evolving adjacency matrices enable successive utilization of MMP layers on these isolated regions in the fine-tuned model.

### v4

# Interpretable A-posteriori Error Indication for Graph Neural Network Surrogate Models

@@ Abstract
Data-driven surrogate modeling has surged in capability in recent years with the emergence of graph neural networks (GNNs), which can operate directly on mesh-based representations of data. The goal of this work is to introduce an interpretability enhancement procedure for GNNs, with application to unstructured mesh-based fluid dynamics modeling. Given a black-box baseline GNN model, the end result is an interpretable GNN model that isolates regions in physical space, corresponding to sub-graphs, that are intrinsically linked to the forecasting task while retaining the predictive capability of the baseline. These structures identified by the interpretable GNNs are adaptively produced in the forward pass and serve as explainable links between the baseline model architecture, the optimization goal, and known problem-specific physics. Additionally, through a regularization procedure, the interpretable GNNs can also be used to identify, during inference, graph nodes that correspond to a majority of the anticipated forecasting error, adding a novel interpretable error-tagging capability to baseline models. Demonstrations are performed using unstructured flow field data sourced from flow over a backward-facing step at high Reynolds numbers, with geometry extrapolations demonstrated for ramp and wall-mounted cube configurations.
@@ 1 Introduction
The goal of this work is to introduce additional, equally critical advantages that can be provided within the GNN framework: (1) the ability to generate interpretable latent spaces (or latent graphs) as a by-product of the network architecture through the addition of a graph sub-sampling module, and (2) the ability to use these interpretable latent graphs to adaptively identify, or tag, sub-graphs that contribute most to the prediction error (i.e., a-posteriori error tagging). To emphasize broad applicability, these goals are achieved here in the context of interpretability enhancements to baseline models: more specifically, when given a pre-trained baseline GNN surrogate model, this work shows how appending an additional graph-based trainable module can not only enhance the interpretability properties of the baseline in various modeling tasks (with focus on mesh-based fluid flow forecasting), but also provide key modeling benefits in the form of error tagging. This strategy is illustrated in Fig. 1.
Figure 1: (Top, blue box) Schematic of baseline GNN model architecture. (Bottom, red box) Schematic of interpretable sub-sampling module. Module is implemented via skip connection between first and last message passing layers in baseline processor block. Interpretability comes from adaptive pooling layer, which generates sub-graphs via sub-sampled nodes that can be visualized as masked fields. Using a regularization strategy, masked fields can then be used to identify regions of forecasting approximation errors during model evaluation. Details provided in Sec. 3.1. (Bottom, green box) Error tagging procedure by means of a budget regularization strategy that encourages prediction errors to be contained within the extracted nodes in sub-graphs (the masked fields). Details provided in Sec. 3.2.
@@
The scope of this previous work [35], however, concerns unstructured flow field reconstruction (i.e., a GNN was used to learn an interpretable identity map). Instead, the objective of this work is to (a) extend this notion of interpretability into a predictive mesh-based modeling setting, with focus on forecasting unstructured turbulent fluid flows, and (b) use such interpretability enhancements to append black-box models with explainable error-tagging capability. As such, given an input graph-based representation of an unstructured flow field, the GNN in this work models the time-evolution of node features. For this forecasting task, the objective is to show how augmenting pre-trained baseline GNNs with trainable adaptive pooling modules results in an interpretable modeling framework for mesh-based predictive modeling. Ultimately, the input to this procedure is an arbitrary GNN baseline model, and the output is a modified GNN that produces interpretable latent graphs tailored to different modeling objectives, including error-tagging capability in the context of data-based forecasting. Although the application focus here is mesh-based fluid flow modeling, the framework is expected to extend into other applications that leverage graph neural networks. The core elements of this work are as follows:
@@
    Multiscale GNN baseline: A baseline GNN leveraging multiscale message passing (MMP) layers is trained to model the evolution of an unstructured turbulent fluid flow. Datasets to demonstrate the methods introduced in this work are sourced from unstructured fluid dynamics simulations of flow over a backward facing step using OpenFOAM, an open-source CFD library [36]. Reynolds numbers of flow trajectories (based on the step height of the backward-facing step configuration) range from 26000 to 46000.
@@
    Interpretable graph sub-sampling module: The aforementioned graph pooling layer – which relies on a learnable node sub-sampling procedure – is attached to the baseline GNN, creating an augmented GNN architecture with enhanced interpretability properties. Interpretability enhancements are then provided by freezing the parameters of the baseline GNN and ensuring the parameters in the newly introduced module are freely trainable. The result is a so-called interpretable GNN.
@@
    Error tagging: A regularization term, representing a mean-squared error budget, is added to the objective function during the interpretable GNN training process. The minimization of this term ensures that the interpretable GNN tags, during inference, a subset of the graph nodes that are expected to contribute most significantly to the GNN forecasting error.
@@
    Geometry extrapolation: interpretable GNNs – which were trained on the BFS geometry alone – are evaluated on two additional geometries exhibiting different flow separation physics (a ramp and a wall-mounted cube case). The objective is to assess the level of consistency in identified features and error tagging capability by means of inference on unseen configurations.
@@ 2 Methods
Mesh-based data required to train the GNNs are sourced from unstructured computational fluid dynamics (CFD) simulations in a two-dimensional backward-facing step (BFS) configuration using OpenFOAM [37]. Simulations are performed for a range of Reynolds numbers (Re) based on the BFS step height to populate training and testing sets. Specifically, 10 training set trajectories were sampled in the equi-spaced Re range of [26214, 44569], and 10 testing set trajectories in the Re range of [27233, 45589] (the reader is referred to Appendix B for a description of corresponding friction Reynolds numbers). Each trajectory is comprised of $419$ two-component velocity snapshots separated by $\Delta t=10^{-4}s$.
@@
    Absence of pressure: The Navier-Stokes equations, which are the governing equations for fluid flow, identify the role of local pressure gradients on the evolution of fluid velocity. The GNN in Eq. 2, however, operates only on velocity field data, and does not explicitly include pressure or its gradients in the evaluation of dynamics. Instead, the model is tasked to implicitly recover the effect of pressure on the flow through data observations. This is consistent with real-world scenarios in which time-resolved velocity data is accessible (e.g., through optical flow [39] or laser-based imaging [40]), but pressure field data (particularly pressure gradient data) is inaccessible or sparse. In the context of incompressible flows (the application considered here), it is possible that explicitly including some pressure contribution, such as a pressure gradient as input, could help with predictions. However, further investigation to this end is required, and such modifications may require either calling a conventional pressure solve or developing some kind of surrogate for the pressure solve (as done in Ref. [41], for example) during inference, essentially making the method physics-informed and perhaps even tied down to an incompressible flow assumption. The intent of this work is to develop a purely data-based approach to mitigate any such assumptions in the model form or objective function.
@@ 3 Results
3.1 Enhancing GNNs for Interpretability
It is assumed that a black-box graph neural network optimized for the above described forecasting task – termed the baseline GNN, ${\cal G}_{B}$ – is available. With ${\cal G}_{B}$ as the starting point, the interpretability enhancement procedure (to be described and demonstrated in this section) produces a new GNN – termed the interpretable GNN, ${{\cal G}_{I}}$ – that is both physically interpretable and optimized for the same modeling task as the baseline ${\cal G}_{B}$. These interpretability properties are embedded into the baseline through architectural modifications; more specifically, ${{\cal G}_{I}}$ is created by appending ${\cal G}_{B}$ with an interpretable sub-sampling module that does not modify the baseline parameters. The module provides interpretability through a learnable pooling operation that adaptively isolates a subset of nodes in the context of the mesh-based forecasting task (the objective function in Eq. 3). It will be shown how these nodes can be readily visualized to extract coherent physical features most relevant to this forecasting goal, all while retaining the architectural structure and predictive properties of the baseline model. This approach is illustrated in Fig. 1.
@@ 3.1.1 Baseline GNN
Message passing layers can take various forms and are the backbone of all GNN architectures, since they directly leverage the graph connectivity to learn complex functions conditioned on the arrangement of graph nodes in neighborhoods [12]. The processor in the baseline architecture in Fig. 1 is a general description, in that the interpretability enhancement procedure is not tied down to the type of MP layer used. Here, all MP layers correspond to multiscale message passing (MMP) layers, which leverage a series of coarse grids in the message passing operation to improve neighborhood aggregation efficiency [16, 35, 15]. Details on the MMP layer used here can be found in Appendix C.
3.1.2 Interpretable Sub-sampling Module
The interpretable sub-sampling module is shown in Fig. 1 (bottom), and contains a separate set of layers designed to enhance interpretability properties of the black-box baseline. The module is implemented through a skip connection between the first and last MP layers in the baseline processor block, such that module parameters are informed of the action of baseline neighborhood aggregation functions during parameter optimization. There are three components to the module: an adaptive graph pooling layer, an additional MMP layer that acts on the pooled graph representation, and a parameter-free un-pooling layer.
@@ 3.1.2 Interpretable Sub-sampling Module
As illustrated in Fig. 1, the sub-graph produced by the pooling layer can be readily visualized in a mesh-based setting, since the extracted nodes correspond to locations in physical space. If the predictive capability of the baseline is withheld when optimizing the parameters in the interpretable module, nodes identified in the sub-graphs serve as an accessible link to the forecasting task. The discussion below shows how adaptive node sub-sampling provided by the pooling operation identifies physically coherent artifacts in the mesh-based domain, thereby connecting the data-based modeling task (the goal of the baseline) with application-oriented features of interest (the goal of the interpretable sub-sampling module).
3.1.3 Interpretability Enhancement Procedure
The interpretability enhancement procedure is accomplished by first initializing the ${{\cal G}_{I}}$ model parameters directly from those in the baseline ${\cal G}_{B}$, which is assumed to already have been trained. Then, the interpretable module is appended to the ${{\cal G}_{I}}$ architecture as per Fig. 1.
A training step for ${{\cal G}_{I}}$ is then carried out using the same objective function used to train the baseline ${\cal G}_{B}$. During training, the subset of parameters in ${{\cal G}_{I}}$ corresponding to the baseline is frozen, and the parameters corresponding to the interpretable module are freely trainable. This ensures that the interpretable GNN parameters include the baseline parameters as a subset, isolating optimization focus to the interpretable module parameters during training. The interpretability enhancement procedure is shown in Fig. 3, which displays objective function optimization histories for interpretable GNNs trained with different values of node reduction factor $RF$.
Figure 3: Mean-squared error objective (Eq. 3) versus training epochs during interpretability enhancement phase.
The training histories show two key trends: (1) the baseline objective history is frozen, as the baseline model ${\cal G}_{B}$ is not being modified, and (2) the interpretable models achieve lower converged mean-squared errors in single-step GNN predictions than the baseline. The latter trend is consistent with the fact that additional parameters are introduced in the interpretability enhancement procedure (see Fig. 1). Due to the fact that a message passing layer acts on the Top-K-identified subgraphs, the converged training errors increase with decreasing RF. Higher RF translates to sub-graphs occupying a smaller percentage of original graph nodes, which in turn limits the scope of the additional message passing layer.
@@ 3.1.4 Masked Fields
The interpretability enhancement is illustrated in Fig. 4, which contrasts the single-step prediction workflow of the baseline GNN with that of a series of interpretable GNNs, each having incrementally higher values of RF. The input to all models is the same instantaneous mesh-based flowfield ${\bf X}_{m}$ sourced from a testing set trajectory.
Figure 4 shows how the interpretability enhancement procedure adds masked fields to the model evaluation output. Simply put, a masked field is generated by visualizing the sub-graph node indices produced by the Top-K pooling operation. Since these truncated nodes coincide with physical space locations consistent with the underlying mesh, the “active” nodes selected during pooling can be extracted and visualized as a node-based categorical quantity (1 if the node was identified and 0 otherwise).
The features contained in these fields allow one to interpret the relationship between the input flow field ${\bf X}_{m}$ and the model-derived forecast $\widetilde{\bf X}_{m+1}$. In other words, through visualization of the masked fields, the interpretable models allow one to access regions in physical space intrinsically linked to the modeling task (single-step prediction), thereby making the model behavior directly accessible from the perspective of the objective function used during training. The masked fields in Fig. 4 show how this added capability translates to interpretable prediction unavailable in the baseline, as the masked fields interestingly isolate nontrivial, but spatially coherent, clusters of nodes (often disjoint in physical space) within the GNN forward pass. Such features can then be connected to the problem physics in an expert-guided analysis phase.
Figure 4: Comparison of GNN prediction workflows between baseline and interpretable GNNs. interpretable workflow includes masked fields generated in forward pass. Input flowfield (top) sourced from Re=27,233 trajectory of the testing set. Bottom-most plots show baseline absolute error fields ($\widetilde{\bf X}_{m+1}-{\bf X}_{m+1}$) for streamwise (left) and vertical (right) velocity components.
@@
Figure 5: Single-step prediction outputs generated by interpretable GNN (RF=16) using input fields at time-ordered instants in the Re=27233 trajectory. Each frame shows predicted y-component of velocity (top) alongside masked field (bottom), with inlays showing subgraph representations.
Since the input velocities stored on the nodes are time-evolving, and the Top-K projection operation is conditioned on these input values, the regions identified in the masked fields are necessarily time-evolving as well. Consequently, the masked fields identify coherent structures that evolve in accordance with the unsteady features of the flow – in the BFS configuration, these features correspond to emerging and downstream shedding of recirculation zones, which in turn lead to a shedding cycle that affects the location at which the flow reattaches to the bottom wall. This time-dependent quality is highlighted in Fig. 5, which shows a set of time-ordered single-step predictions paired with corresponding masked fields during one BFS shedding cycle. It should be noted that the adaptive pooling operation induces a sub-graph from which the masked field is derived; since the identified regions in the masked fields are time-evolving, the adjacency matrix that characterizes this sub-graph is time-evolving as well, and remains coincident with the fixed adjacency matrix of the original mesh-based graph. The sub-graphs are provided as inlaid plots in Fig. 5; these evolving adjacency matrices enable successive utilization of MMP layers on these isolated regions in the interpretable model.